If the refractive index of glass is 1.5, what is the critical angle at the air-glass
interface?
1
Expert's answer
2019-09-24T09:13:03-0400

"\alpha = \arcsin\frac{n_{air}}{n_{glass}} = \arcsin\frac{1}{1.5} \approx 41.8 \degree."